# Temporal Valued Constraint Satisfaction Problems

### v1

@@ 5.3. Classification
Proof.
Let ${\mathfrak{A}}^{\prime}:=(\mathbb{Q};\langle{\mathfrak{A}}\rangle_{0}^{\infty})$. By Theorem 2.25, $\operatorname{VCSP}({\mathfrak{A}})$ is in NP. By Proposition 3.2, ${\mathfrak{A}}$ pp-constructs $K_{3}$ if and only if ${\mathfrak{A}}^{\prime}$ pp-constructs $K_{3}$ and in this case, $\operatorname{VCSP}({\mathfrak{A}})$ is NP-complete by Lemma 2.16. Hence, it follows from Theorem 5.9 applied on ${\mathfrak{A}}^{\prime}$ that either (1) holds or $\operatorname{Pol}(\mathbb{Q};\langle{\mathfrak{A}}\rangle_{0}^{\infty})$ contains a pwnu operation. Hence, if $\operatorname{Pol}(\mathbb{Q};\langle{\mathfrak{A}}\rangle_{0}^{\infty})$ contains a pwnu operation, then ${\mathfrak{A}}$ does not pp-construct $K_{3}$. By Proposition 5.10 and Theorem 4.1, $\operatorname{Betw}$, $\operatorname{Cycl}$, $\operatorname{Sep}$, $T_{3}$, $-T_{3}$, $\operatorname{Dis}\notin\langle{\mathfrak{A}}\rangle$ and therefore, item (2) from Corollary 5.26 applies and $\operatorname{VCSP}({\mathfrak{A}})$ is in P. ∎
Conjecture 9.3 in [15] states that, under some structural assumptions on ${\mathfrak{A}}$, $\operatorname{VCSP}({\mathfrak{A}})$ is in P whenever ${\mathfrak{A}}$ does not pp-construct $K_{3}$ (and is NP-hard otherwise) [^2] The original formulation uses the structure ($(\{0,1\};\operatorname{OIT})$, but it well-known that this structure pp-constructs $K_{3}$ and vice versa [3].. All temporal structures satisfy the assumptions of the conjecture and hence Corollary 5.27 confirms the conjecture for the class of temporal VCSPs.
6. Concluding Remarks
We proved a complexity dichotomy for temporal VCSPs: $\operatorname{VCSP}({\mathfrak{A}})$ is in P or NP-complete for every valued structure ${\mathfrak{A}}$ such that $\operatorname{Aut}(\mathbb{Q};<)\subseteq\operatorname{Aut}({\mathfrak{A}})$. Moreover, we showed that the meta-problem of deciding whether $\operatorname{VCSP}({\mathfrak{A}})$ is in P or NP-complete for a given ${\mathfrak{A}}$ is decidable. As a side product of our proof, we obtain that the complexity of every such VCSP is captured by the classical relations that it can express, in other words, by the CSPs that are encoded in this VCSP. Our results confirm [15, Conjecture 9.3] for all temporal valued structures.
The proof of our decidability result (Proposition 5.25) is based on the distinction of two cases depending on whether P=NP. The typical results on decidability of these meta-problems in the theory of (V)CSPs are rather formulated by deciding the algebraic conditions that imply the respective complexities, more concretely, deciding the presence of certain (fractional) polymorphisms. This can often be checked by the naive approach, as long as the signature of the structure is finite. However, if we wanted to do so in our case, we would have to check for polymorphisms of the structure $(\mathbb{Q};\langle{\mathfrak{A}}\rangle_{0}^{\infty})$, which has an infinite signature by definition. This motivates the following question.
Question 6.1.
Let ${\mathfrak{A}}$ be a valued structure with a finite signature such that $\operatorname{Aut}(\mathbb{Q};<)\subseteq\operatorname{Aut}({\mathfrak{A}})$. Given a first-order definition of a valued structure ${\mathfrak{A}}$ with a finite signature in $(\mathbb{Q};<)$, is it decidable whether ${\mathfrak{A}}$ pp-constructs $K_{3}$, equivalently, whether item (2) in Corollary 5.26 holds?
@@ 6. Concluding Remarks
A positive answer to the second question in Question 6.2 would confirm [15, Conjecture 9.3] for valued structures preserved by all automorphisms of the countable random graph.

### v2

@@ 5.4 Classification
Proof 20
Let ${\mathfrak{A}}^{\prime}:=(\mathbb{Q};\langle{\mathfrak{A}}\rangle_{0}^{\infty})$. By Theorem 21, $\operatorname{VCSP}({\mathfrak{A}})$ is in NP. By Proposition 25, $\mathfrak{A}$} $p$-constructs $K_{3}$ if and only if $\mathfrak{A}$$^{\prime}$ $p$-constructs $K_{3}$ and in this case, $\operatorname{VCSP}({\mathfrak{A}})$ is NP-complete by Lemma 12. Hence, it follows from Theorem 45 applied to $\mathfrak{A}$$^{\prime}$ that either item 1 holds or $\operatorname{Pol}(\mathbb{Q};\langle{\mathfrak{A}}\rangle_{0}^{\infty})$ contains a pwnu operation. Hence, if $\operatorname{Pol}(\mathbb{Q};\langle{\mathfrak{A}}\rangle_{0}^{\infty})$ contains a pwnu operation, then $\mathfrak{A}$} does not pp-construct $K_{3}$. By Proposition 46 and Theorem 27, $\operatorname{Betw}$, $\operatorname{Cycl}$, $\operatorname{Sep}$, $T_{3}$, $T_{3}$, $\operatorname{Dis}\notin\langle{\mathfrak{A}}\rangle$ and therefore, item 2 from Corollary 57 applies and $\operatorname{VCSP}({\mathfrak{A}})$ is in P. $\qed$
Conjecture 9.3 in [16] states that, under some structural assumptions on $\mathfrak{A}$}, $\operatorname{VCSP}({\mathfrak{A}})$ is in P whenever $\mathfrak{A}$} does not pp-construct $K_{3}$ (and is NP-hard otherwise) [^2] The original formulation uses the structure $(\{0,1\};\operatorname{OIT})$, but it is well-known that this structure pp-constructs $K_{3}$ and vice versa [17].. All temporal valued structures satisfy the assumptions of the conjecture and hence Corollary 58 confirms the conjecture for the class of temporal VCSPs.
6 Future work
@@
A positive answer to the second question in Question 59 would confirm [16, Conjecture 9.3] for valued structures preserved by all automorphisms of the countable random graph.
Acknowledgements
Manuel Bodirsky and Žaneta Semanišinová have been funded by the DFG (Project FinHom, Grant 467967530) and by the European Research Council (Project POCOCOP, ERC Synergy Grant 101071674). Views and opinions expressed are however those of the authors only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them.
Édouard Bonnet was supported by the ANR project TWIN-WIDTH (ANR-21-CE48-0014-01).
Žaneta Semanišinová was funded in whole or in part by the Austrian Science Fund (FWF)10.55776/ESP6949724.